My Not 1/3 points 1 Previous Answers OSUniPhys1 10.3 P053 The angular velocity of a flywheel with radius 2.0 m varies according to ω(t) 6.0t, pot adr) and at(t) from t-0 to 1.0 s for r-2.0 m, a (m/s a (m/s) 70 60 60 50 a(t) alt) 40 30 20 10 40 30 ntr) in t (s) t (s) 0.2 0.4 0.6 0.8 1.0 O 0.2 0.4 0.6 08 1.0 a (m/s) a (m/s) 70 atn 70 60 ate)
media%2Fa44%2Fa441b233-cfc0-4d95-981e-35
Answer #1

\omega  = 6t

Angular acceleration \alpha = d\omega/dt = 6

Tangential acceleration is given by at = r\alpha = 2*6 = 12 m/sec2 ...... constant

Centripetal acceleration is given by ac = r\omega2

ac = 2*(6t)2 = 72t2

Graphs of ac :

Graph of at :

From above graph correct answer is (a)

From fig(a) at >>ac when time is near to zero and

ac >>at when time goes to infinity or very big